OpenStudy (kanika):

How v will find d value of tan 22.5?

imqwerty (imqwerty):

tan(2x)=[2tanx]/[1-tan^2(x)] Try using this

zepdrix (zepdrix):

Tangent Half-Angle Identity should help us here:\[\large\rm \tan\left(\frac{x}{2}\right)=\frac{1-\cos x}{\sin x}\]Rewrite your 22.5 as half of 45 degrees.

zepdrix (zepdrix):

\[\large\rm \tan(22.5)=\tan\left(\frac{45}{2}\right)=\frac{1-\cos45}{\sin45}\]

OpenStudy (mathstudent55):

This question depends on what your conditions are. If the are are no restrictions, use a calculator to find an approximate answer. If you need an exact answer, use the half-angle identity. The sine and cosine for 45 degrees are well known exact values.